You have credit card debt of $4,000 at 23% APR compounded monthly. If you charge no more purchases to the card and make monthly payments of $375, how many months will it take you to payoff your debt?
Answer #1

PV =

[1-(4+5)-) [1-(1+r)-n LT P= Periodic Payment r=rate per period n = number of periods

r = 23%/12 = 1.9167% (per month)

P = $375

PV = $4000

4000 = 375 * 1 4000 = 375 *1-11 + 0.019167) -n 0019167

0.204444 = 1 - (1.019167)-n

(1.019167)-n = 0.795556

Taking log of both sides,

-n * LN(1.019167) = LN(0.795556)

-n * 0.018984 = -0.22871

n = 12.05 months